# Yet Another Mirage of Breaking MIRAGE: Debunking Occupancy-based Side-Channel Attacks on Fully Associative Randomized Caches

### v2

@@ I Introduction
First, we tried to reproduce the results of the RCO paper. Using the unmodified artifact released with the RCO paper [^1] We use the the code artifact released with RCO’s paper [5] - https://zenodo.org/records/14869981 which corresponds to the latest release on their Github repository, v2.0 (commit: fc07ea6), at the time of writing., we attempted to reproduce their AES key recovery results, specifically the guessing entropy for an unknown victim AES key reported in Figure 10 of the RCO paper. [5]. Surprisingly, we were unable to reproduce the entropy degradation for MIRAGE as reported by them. Figure 1 shows the guessing entropy of an unknown AES key (higher is better), as reported by RCO for MIRAGE and other randomized cache designs, and our own reproduction using RCO’s artifact.
As shown in Figure 1, the RCO paper reports the guessing entropy to decrease for MIRAGE as the number of AES encryptions increases. However, our reproduction [^2] Our reproduction code and scripts for all the figures in this paper is open-sourced at https://github.com/sith-lab/yet-another-mirage-of-breaking-mirage. of their result shows that MIRAGE has high entropy (above 90%) even after thousand AES encryptions, similar to other randomized cache designs such as CEASER-S, SassCache, and ScatterCache.
Figure 1: Guessing Entropy (GE) for an unknown AES key, as the number of AES encryptions increases, for our reproduction of MIRAGE from RCO’s artifact compared to results from the RCO paper for MIRAGE, CEASER-S, SassCache, ScatterCache. In our reproduction, MIRAGE has GE of more than 90%, showing no leakage, departing from the RCO paper.
Next, given that the entropy results in RCO could not be reproduced, we examine whether the stated root cause of the reported leakage is valid. RCO attributes the leakage to the accesses to the AES S-Box in the last round of AES encryption and its impact on cache occupancy, and its correlation with an attacker’s access time for its own cached array. By creating access time templates for each possible key-byte value of a profiled key, the attacker matches the observed timings for an unknown victim key against the templates to recover the key. However, this fails to explain how the attack works around the noise introduced by random global evictions in MIRAGE.
@@
Effect of Random Global Evictions. MIRAGE randomly evicts an existing LLC line upon every new insertion. This means that even repeated AES encryptions of the same plaintext and key would produce different cache occupancy patterns, due to the different eviction patterns, unrelated to the key. Moreover, an attacker accessing its own array to estimate the cache occupancy of the victim, would see further noise due to its own addresses getting randomly evicted differently on each run. Thus, it seems difficult for such an attacker to even fingerprint a particular key, let alone recover the entire key byte by byte under random global evictions. In investigating this, we uncover a modeling bug in RCO’s evaluation.
@@
Modeling Flaw in RCO. RCO’s simulations initialize MIRAGE’s global eviction RNG with a static seed before each AES encryption, causing a fixed sequence of evictions on each AES encryption, unlike real hardware.
@@
RCO runs each AES encryption in a separate simulation starting from a clean cache state. In this setup, the final cache occupancy is determined by both the victim’s accesses and the global eviction pattern. However, with a fixed RNG seed, the eviction pattern never changes, making occupancy entirely a function of the victim’s access pattern and spuriously creating correlations with the key. If we randomize the seed for each run (e.g., using std::random_device [^3] Note that on some systems without a hardware-based entropy source, std::random_device can also result in a deterministic output [7]. Please validate on your system if the source of randomness you use is truly random. to seed the RNG), the eviction pattern changes on each run, and the correlation between the key and attacker timings disappears, as expected in MIRAGE. This models a realistic setting, where the attacker cannot reset or guess the state of the RNG for the evictions.
@@
Results with Correct Modeling. In evaluations of an unknown victim key’s guessing entropy (GE), after randomizing the RNG seed for global evictions in MIRAGE, the GE remains high (>>90%), highlighting that AES key leakage on MIRAGE via occupancy attacks is infeasible.
@@
    We show that the AES key guessing entropy remains high in MIRAGE, in a reproduction of RCO’s artifact, contradicting the paper’s claims.
@@
    We identify that RCO’s fixed RNG seeds for global evictions in MIRAGE create artificial correlations, and randomizing the seed, as in real hardware, eliminates any observed correlation and leakage.
@@ II-B Cache Occupancy Attacks on Mirage
Cache occupancy attacks measure changes in the overall occupancy of a shared cache, rather than targeting specific sets as in set-conflict attacks like Prime+Probe. Because they exploit aggregate cache usage, all randomized caches without explicit cache partitioning, such as CEASER, ScatterCache, and Mirage in principle, leak some information via cache occupancy. In fact, MIRAGE’s threat model explicitly claims to not protect against such occupancy-based channels. While covert channels between two colluding processes can be naively constructed by modulating the cache occupancy, the RCO [5] paper claims a stronger side-channel: recovering AES keys through modulation of cache occupancy.
@@ III Analyzing Claims of Occupancy-Based Side-Channel Attacks on MIRAGE
Using the authors’ publicly released artifact, we attempted to reproduce the results in Figure 10 of the RCO paper. However, contrary to their claims, we found that the guessing entropy for an unknown AES key remained high and did not drop even after thousands of encryptions in our reproduction, as shown in Figure 1. Therefore, we analyze the claimed root cause of the attack: that the execution time for an attacker accessing its own array depends on the cache occupancy of the AES encryption, which is a function of the AES key.
@@ III-A Pitfall-1: Fixed Sequence of Evictions Modeled by RCO
RCO Attack Root Cause. The RCO [5] paper claims that the T-Table implementation of AES running on MIRAGE [1] can leak the secret key through a cache occupancy attack. The root cause of this leakage as mentioned by RCO is that the sequence of T-Table accesses in the last round of AES encryption, which depends on the secret key, can impact the cache occupancy in MIRAGE. Therefore, the execution time for an attacker to access a large cached array, parts of which may have been evicted by the AES encryption, can leak the cache occupancy, and therefore the secret key.
@@
Our Reproduction of RCO’s Root Cause. Since we are unable to reproduce the exact Guessing Entropy results from RCO (see Figure 1) , we try to reproduce their root cause, the template with correlation between profiled and victim keys. Figure 2(a) shows a heatmap visualizing the template built by the attacker, where each entry represents one of the 256 possible T-table entries (arranged as a $\times16$ matrix) and the cell color encodes the attacker’s average access time, when that entry is accessed in the last round of AES by the victim. If the profiled-key heatmap (attacker’s template) closely matches the heatmap with the guessed victim-key, the guess is likely to be the correct key. For simplicity, instead of $T_{n}=\text{SBOX-INV}(K\oplus CT)$, we use $T_{n}=K\oplus CT$ for our bins, since SBOX-INV is just a lookup table, and visualize a single heatmap, averaging the heatmaps of key bytes 0 to 15.
@@
As shown in Figure 2(a), when we generate the heatmaps by using the RCO artifact [5] there is a clear correlation between the templates of the profiled key and the guessed victim key. This correlation can allow the attacker to leak the victim key since the templates will be correlated only when the guessed key is actually the correct victim key.
RCO’s Bug: Fixed Sequence of Global Evictions. To understand why this correlation exists, we measured the attacker access times, after victim AES encryptions, for four sample plaintext-ciphertexts pairs chosen at random and repeated 100 times. Each of these encryptions has a distinct AES T-Table access sequence in the last round. Figure 3(a) shows the histogram of the access-times measured for these four encryptions: each of these (CT 1, 2, 3) has a different access time, although CT 3 and 4 overlap with the same access time. This explains the correlations of access-times with AES T-table access sequences. Surprisingly, all 100 repetitions yield the exact same measurement for the attacker access time, although global evictions during the attacker accesses are supposed to be random. This allowed us to identify the bug in the RCO implementation, that the Global Evictions used a random-number generator (RNG) seeded with a static seed (42). Hence, in each AES execution, the same sequence of cache indices is selected for global eviction, making the operation of MIRAGE quite deterministic in each AES encryption, which is quite unrealistic in real implementations, where the attacker has no way to reset the RNG to a fixed state each time.
@@
Our Fix: Accurate Modeling with Random Seed. The cache occupancy ($O$) is a function of both the Victim accesses ($V$) and the Global Eviction decisions ($GE$), i.e., $O\approx f(V,GE)$. RCO incorrectly assumes a fixed sequence of GE for each encryption, making the fingerprinting of $V$ using $O$ measurements possible. Fixing the bug in RCO, by changing the fixed seed for global evictions to a randomly chosen seed for each AES encryption, representative of real systems, where the Global Eviction decisions would be performed by a RNG whose state cannot be reset by the attacker, we see that the $CO$ is now strongly impacted by the $GE$ in addition to the $V$. As shown in Figure 3(b), with a random seed for each encryption, the attacker’s access times show random variations of the order of 100,000 cycles, due to the global evictions unpredictably evicting the attacker’s own lines during its measurement phase, that overwhelm minor occupancy differences causing by the victim which varied timings by a few 1000 cycles in Figure 3(a).
@@
After our fix, when we replace the fixed seed with a random seed for each AES encryption in Figure 2(b), the correlation between profiled and victim keys disappears, making the heatmaps virtually unrelated and eliminating the signal required for key recovery. This realistic modeling of MIRAGE’s global evictions removes the attack’s timing signal, preventing AES key leakage on MIRAGE via occupancy attacks.
@@ III-B Pitfall-2: Unrealistic L1 Cache Configuration
The RCO paper also claims to use a 512 kilobyte L1 cache, as per Section 7.1 of their paper [5]. However, their code artifact [^4] We refer to the code artifact version 2.0 released by the authors on Zenodo - https://zenodo.org/records/14869981 uses a 512 byte L1 cache by default which is not representative of real systems which use at least a 64KB L1 Cache. Modeling a smaller L1 cache, such as 512 byte L1 cache, can inflate the L1 cache misses and LLC accesses, compared to a 64KB L1 Cache which can have hits for all T-table accesses after the first round. Thus, modeling a 512 byte L1 cache can overestimate the attack. We confirmed that after updating the initialized seed to be random, regardless of the L1 cache size being 512-byte like in the RCO artifact or a more realistic 64KB, there is a lack of correlation between the heatmaps of the profiled and victim key templates, similar to Figure 2 (b), indicating that AES key leakage is impractical.
@@ III-C Guessing AES Key after Fixing RCO’s Modeling Issues
We evaluate the Guessing Entropy (GE) for a victim’s AES key, using the template for a profiled AES key similar to RCO, using the formula: $GE=\sum_{i=0}^{15}log_{2}(R_{i})$, where $R_{i}$ is the rank of the correct guess for key byte $i$. Figure 4 shows the guessing entropy (GE) for Mirage from the RCO paper, our reproduction based on their artifact, and after our fixes of RCO’s modeling issues (after using a random seed for initializing the RNG for global evictions). In our reproductions (both before and after our fixes), we continue to see Mirage with high GE, demonstrating that it is resilient to brute-force key guessing attacks on AES.
Figure 4: Guessing Entropy (GE) for an unknown AES key, after our fixes, as the number of AES encryptions increases. MIRAGE continues to have a high GE of more than 90%, showing no leakage, with Random Seed used to initialize the RNG for Global Evictions, with both an unrealistic 512B L1 Cache (like RCO’s artifact) and a realistic 64KB L1 Cache.

### v3

@@ I Introduction
First, we attempted to reproduce the AES key recovery results reported in the RCO paper using their artifact [^1] We use the the code artifact with RCO’s paper [5] - https://zenodo.org/records/14869981 (Github release v2.0, commit: fc07ea6)., specifically the guessing entropy (GE) for an unknown victim key (Figure 10 in the RCO paper [5]). Using the RCO artifact, we initially could not reproduce their results.
Bug-1: GE Analysis Bug in RCO. Figure 1 compares the guessing entropy (GE) for an AES key (higher is better) for MIRAGE using RCO’s raw data (RCO – red line) and our reproduced data (Reproduced – blue line), along with other randomized caches. In the original analysis (Figure 1a), RCO reports GE dropping below 80 after just 300 AES traces. In contrast, our reproduced data shows the same GE reduction occurs after 1400 traces.
[…]
Figure 1: Guessing Entropy (GE) for an unknown AES key, as the number of AES encryption traces increases. (a) Original analysis from RCO shows GE reduction for MIRAGE after 300 traces. (b) After fixing the analysis bug, GE reduction occurs after 1400 traces, consistent with our reproduction.
Next, we examine the validity of this leakage.
@@
Bug-2: Global Eviction Modeling Bug in RCO. RCO attributes AES key leakage to last-round S-Box accesses in the AES encryption influencing cache occupancy. The attacker observes this by measuring access times to its own cached array, building templates for each key-byte value using a profiled key, and matching these templates against timings for an unknown key to recover it. However, this explanation ignores MIRAGE’s random global evictions, which introduce noise in cache occupancy. MIRAGE randomly evicts an existing LLC line upon every new insertion. Hence, repeated AES encryptions of even the same plaintext and key evict different addresses and produce different occupancy (unrelated to the key), and the attacker’s own accesses also evict its addresses randomly on each run, introducing further noise. Thus, an attacker would even find fingerprinting a key difficult (Figure 3), let alone leak the entire key byte-wise. Investigating this, we uncover a second bug in RCO’s evaluation:
@@
Modeling Flaw in RCO. RCO’s simulations initialize the RNG used by MIRAGE’s evictions with a static seed before each AES encryption, causing a fixed sequence of evictions on each AES run, not practical in a real attack.
RCO’s simulations seed MIRAGE’s RNG used for evictions with a fixed value (42) before each AES encryption, restarting each AES run from an identical state (see Appendix B). This produces a fixed eviction sequence on each AES encryption, not representative of MIRAGE. In a realistic modeling of MIRAGE with random global evictions, the final cache occupancy (O) after each AES encryption should be determined by both the victim’s accesses (V) and the global random eviction sequence (R), $O\approx f(R,V)$. However, as RCO uses a constant RNG seed, producing a fixed eviction sequence in each run, the cache occupancy becomes a function of just victim accesses, $O\approx f(V)$, artificially creating correlations with the key.
@@
To correctly model MIRAGE while restarting the simulation for each AES run, we randomize the eviction RNG seed for each AES run (e.g., using std::random_device [^3] Note that on some systems without a hardware-based entropy source, std::random_device can also result in a deterministic output [7]. Please validate that the source of randomness you use is truly random on your system. to seed the RNG). This models a realistic setting, where the attacker cannot reset the RNG to a constant state on each AES run. With this correction, correlations between attacker timings and keys disappear, and the victim AES key’s guessing entropy remains high (above 90%), as shown in Figure 4.
@@
Results with Correct Modeling. After randomizing the RNG seed for MIRAGE in each AES run, the victim key’s guessing entropy (GE) remains high (>>90%), see Figure 4. Thus, AES key leakage on MIRAGE is infeasible.
@@
    We show that the AES key guessing entropy remains high in MIRAGE, contradicting the RCO paper’s claims.
@@
    We identify that RCO’s usage of a fixed RNG seed for MIRAGE’s evictions creates artificial correlations; randomizing the seed eliminates any observed leakage.
@@ II-B Cache Occupancy Attacks on Mirage
Cache occupancy attacks measure changes in the overall occupancy of a shared cache, rather than targeting specific sets as in set-conflict attacks like Prime+Probe. Because they exploit aggregate cache usage, all randomized caches without explicit cache partitioning, such as CEASER, ScatterCache, and Mirage in principle, leak some information via cache occupancy. In fact, MIRAGE’s threat model explicitly claims to not protect against such occupancy-based channels. While covert channels between two colluding processes can be naively constructed by modulating the cache occupancy, the RCO [5] paper claims the existence of a stronger side-channel on MIRAGE: leaking AES keys via cache occupancy effects.
@@ III Analyzing Claims of Occupancy-Based Side-Channel Attacks on MIRAGE
Using the authors’ publicly released artifact, we first reproduced the results in Figure 10 of the RCO paper, as shown in Figure 1. Once we were able to reproduce the trend, that the guessing entropy for an unknown AES key reduces after 1400 encryptions, we analyze the root cause of the attack: that the execution time for an attacker accessing its own array is a function of the cache occupancy of the AES encryption, which is in turn a function of the victim’s AES key.
@@ III-A Pitfall-1: Fixed Sequence of Evictions Modeled by RCO
RCO Attack Root Cause. The RCO [5] paper claims that the T-Table implementation of AES running on MIRAGE [1] can leak the secret key through a cache occupancy attack. The root cause of the leakage claimed by RCO is that the sequence of T-Table accesses in the last round of AES encryption, which depends on the secret key, can impact the cache occupancy in MIRAGE. Therefore, the execution time for an attacker to access a large cached array, parts of which may have been evicted by the AES encryption, can leak the cache occupancy, and therefore the secret key.
@@
Reproducing RCO’s Root-Cause. To validate RCO’s leakage, we try to reproduce its root cause, the templates with correlation between profiled and victim keys. Figure 2(a) shows a heatmap visualizing the template built by the attacker, where each entry represents one of the 256 possible T-table entries (arranged as a $\times16$ matrix) and the cell color encodes the attacker’s average access time, when that entry is accessed in the last round of AES by the victim. If the profiled-key heatmap (attacker’s template) closely matches the heatmap with the guessed victim-key, the guess is likely to be the correct key. For simplicity, instead of $T_{n}=\text{SBOX-INV}(K\oplus CT)$, we use $T_{n}=K\oplus CT$ for our bins, since SBOX-INV is just a lookup table, and visualize a single heatmap, averaging the heatmaps of key bytes 0 to 15.
@@
As shown in Figure 2(a), when we generate the heatmaps by using the RCO artifact [5] there is a clear correlation between the templates of the profiled key and the guessed victim key. This correlation can allow the attacker to leak the victim key since the templates will be highly correlated when the guessed key is actually the correct victim key.
RCO’s Bug: Fixed Sequence of Global Evictions. To investigate the source of the observed correlation, we measured attacker access times after victim AES encryptions for four randomly chosen plaintext-ciphertext pairs, each repeated 100 times. Each encryption has a distinct last-round AES T-Table access sequence. Figure Figure 3(a) shows the histogram of attacker access times. CT1, CT2 and CT3 differ in access times (although CT3 and CT4 overlap), highlighting that the access times are indeed correlated with T-Table accesses. However, notably, all 100 repetitions of each encryption produce identical access-time measurements, despite global evictions being intended as random.
This reveals a bug in RCO: the global eviction RNG is seeded with a static value (42). As a result, each AES execution follows the same fixed eviction sequence, making MIRAGE’s behavior deterministic. In a real implementation, the attacker cannot reset the RNG between runs, so the deterministic behavior is unrealistic and artificially creates correlations between access times and AES T-Table sequences.
@@
Our Fix: Accurate Modeling with Random Seed. The cache occupancy ($O$) in MIRAGE is a function of both the Victim accesses ($V$) and the Global Eviction decisions ($GE$), i.e., $O\approx f(V,GE)$. RCO incorrectly assumes a fixed sequence of GE for each encryption, making the fingerprinting of $V$ using $O$ measurements possible. We fix this bug in RCO, by changing the fixed seed for global evictions to a randomly chosen seed for each AES encryption simulation, representative of real systems, where the Global Eviction is performed by a RNG whose state cannot be reset by the attacker. After this bug fix, we see that the $O$ is now strongly impacted by the $GE$ in addition to the $V$. As shown in Figure 3(b), with a random seed for each encryption, the attacker’s access times show random variations of the order of 100,000 cycles, due to the global evictions unpredictably evicting the attacker’s own lines during its measurement phase. These overwhelm minor occupancy differences caused by the victim which varied timings by few 1000 cycles in Figure 3(a).
@@
After our fix, when we use a random seed for each AES encryption in Figure 2(b), the correlation between profiled and victim keys disappears, making the heatmaps virtually unrelated, eliminating the signal required for key leakage. This realistic modeling of MIRAGE’s global evictions removes the attack’s timing signals, preventing AES key leakage on MIRAGE via occupancy attacks. We provide the code for this bug fix in Appendix B.
@@ III-B Pitfall-2: Unrealistic L1 Cache Configuration
The RCO paper claims to use a 512 kilobyte L1 cache, as per Section 7.1 of their paper [5]. However, their code artifact [^4] We refer to the code artifact version 2.0 released by the authors on Zenodo - https://zenodo.org/records/14869981 uses a 512 byte L1 cache by default which is not representative of real systems which typically possess atleast 64KB L1 Cache. Modeling a smaller L1 cache, such as 512 byte L1 cache, can inflate the L1 cache misses and LLC accesses, compared to a 64KB L1 Cache which can provide hits for all T-table accesses after the first round. Thus, a 512 byte L1 cache can overestimate the attack success. We confirmed that after updating the initialized seed to be random, regardless of the L1 cache size being 512-byte like in the RCO artifact or a more realistic 64KB, there is a lack of correlation between the heatmaps of the profiled and victim key templates, similar to Figure 2 (b), indicating that AES key leakage is impractical in MIRAGE.
@@ III-C Guessing AES Key after Fixing RCO’s Modeling Issues
We evaluate the Guessing Entropy (GE) for a victim’s AES key, using the template for a profiled AES key similar to RCO, using the formula: $GE=\sum_{i=0}^{15}log_{2}(R_{i})$, where $R_{i}$ is the rank of the correct guess for key byte $i$. Figure 4 shows the guessing entropy (GE) for Mirage based on their artifact (reproduced), and with our fixes of RCO’s modeling issues where we use a random seed to initialize the RNG for global evictions (bug fix). After our bug fixes, we see that Mirage has high GE above 90%, even after thousands of AES encryptions, demonstrating that it is resilient to brute-force key guessing attacks on AES when modeled correctly.
Figure 4: Guessing Entropy (GE) for an unknown AES key, after our fixes, as number of AES encryptions increases. MIRAGE has a high GE above 90% even after thousands of AES encryptions, showing no leakage, once the Eviction RNG is initialized with a Random Seed, with both a 512B L1 Cache (like RCO’s artifact) and a more realistic 64KB L1 Cache.
@@ V Acknowledgment
Appendix A Bug-1: Guessing Entropy Analysis Bug in RCO
Bug def develop_tuples( round_timing_tuples, filename, key_position): datalines = open(filename). readlines() index = 0 for data in datalines: if(index > max_traces): continue index = index + 1 .. % process trace for filename in all_files: develop_tuples(..)
Fix def develop_tuples( round_timing_tuples, filename, key_position, traces_used): datalines = open(filename). readlines() for data in datalines: if(traces_used > max_traces): continue traces_used = traces_used + 1 .. % process trace return traces_used traces_used = 0 for filename in all_files: traces_used = develop_tuples(.., traces_used)
Figure 5: Bug in RCO’s Guessing Entropy Code
Figure 5 illustrates the bug in RCO’s Guessing Entropy Analysis code. In the original implementation (left), RCO processes multiple trace files (up to 6 AES trace files) to calculate the guessing entropy, creating tuples for computing guessing entropy. The bug arises because the check against max_traces is applied locally per trace file, rather than globally across all trace files. As a result, the code uses ends up using $\texttt{6}\times\texttt{max\_traces}$, instead of the intended max_traces, causing the number of AES traces needed for a given GE to be under-reported by a factor of six.
In our bug fix (right), we introduce a traces_used variable that tracks the total number of traces used across all files. This ensures that the max_traces limit is enforced globally correcting this counting error.
Our reproduction stored all the traces in a single file and thus was not affected by this bug. The discrepancy between the reproduced results and RCO’s reported results helped us identify this issue. The code for this bug fix is available here.
Appendix B Bug-2: Global Eviction RNG Seed Bug in RCO
Bug VwayTags(const Params *p) :BaseTags(p),.. replacementPolicy( p->replacement_policy) mt_rand(42) // constant seed { .. }
Fix VwayTags(const Params *p) :BaseTags(p),.. replacementPolicy( p->replacement_policy), { std::random_device rd; mt_rand.seed(rd()); // random seed }
Figure 6: Bug in RCO’s modeling of MIRAGE’s Evictions
Figure 6 illustrates the bug in RCO’s modeling of MIRAGE’s evictions. In the original code (left), the RNG used for global tag evictions is initialized with a fixed seed (42). Due to RCO’s attack methodology of restarting the simulation for each AES encryption, this causes every AES simulation to artificially produce the same eviction sequence. The bug fix (right) replaces the fixed seed with a random seed unpredictable to an attacker generated via std::random_device, producing a different global eviction sequence for each AES simulation. This models a realistic behavior of an attack on MIRAGE. Our patch for the bug fix is available at this link.